bill533:

Which phrase describes the pattern? 1, 2, 5, 14, 41,... A) add 1 B) add 9 C) multiply by 2 D) multiply by 2 and add 4 E) multiply by 3 and subtract 1

bill533:

@mhchen

Gdeinward:

ideas?

bill533:

i believe it is C

Gdeinward:

that is incorrect. Try again

bill533:

E) multiply by 3 and subtract 1

Gdeinward:

that is correct. Good job

Elsa213:

We know it's not the first choice because then it'll be `1,2,3,4,5, etc.` It can't be the second choice because it'll be `1, 10, 19, etc.` The third choice is incorrect because 2*2=4 and there is no four. e.e 1*3= 3 When you subtract 1, you get 2. 2*3=6 Subract 1= 5 And so on... The last choice is your answer.

AngeI:

Bill I feel like you just guessed until you got it right. So you can test out which one it is For example andnevermindElsadidit

bill533:

No i checked over and i did calculations 41*3=123 123-1=122

Gdeinward:

Bill normally does, hence why I did not bother to explain.